A boat can travel 120 miles upstream against the current in the same amount of time it can travel 180 miles downstream with the current. If the boat's average speed in still water is 20 miles per hour, find the speed of the current.
Answer by mananth(16949)

A boat can travel 120 miles upstream against the current in the same amount of time it can travel 180 miles downstream with the current. If the boat's average speed in still water is 20 miles per hour, find the speed of the current.
..
let current speed be x
..
upstream speed =20-x
distance = 120 miles
time = 120/(20-x)
..
down stream speed = x+20
distance = 180 miles
time = 180/(x+20)
...
times are equal
120/(20-x)=180/(x+20)
120(x+20)=180(20-x)
120x+2400=3600-180x
120x+180x=3600-2400
300x=1200
x= 4 mph the speed of current.
...
m.ananth@hotmail.ca